I have an quadratic equation that has to do with graphing and its due tomorrow night (4/16/2010). P=7t^2+8000. 8000 represents the population (p0), t is the time meaning years. t=o (2010), t=1 (2011), t=2 (2012), t=3 (2013), and t=6 (2016). The time (year) is already solved for me but I dont know how to figure out the solving of the population for the years that I have shown nor graph it all onto an graph. Can you please help me.
Thank you,
tpickens

Answer by nerdybill(7384)

P=7t^2+8000. 8000 represents the population (p0), t is the time meaning years. t=o (2010), t=1 (2011), t=2 (2012), t=3 (2013), and t=6 (2016). The time (year) is already solved for me but I dont know how to figure out the solving of the population for the years that I have shown nor graph it all onto an graph.
.
t=o (2010)
P(t)=7t^2+8000
P(0)=7(0)^2+8000
P(0)=8000
.
t=1 (2011)
P(t)=7t^2+8000
P(1)=7(1)^2+8000
P(1)=7+8000
P(1)=8007
.
t=2 (2012)
P(t)=7t^2+8000
P(2)=7(2)^2+8000
P(2)=28+8000
P(2)=8028
.
t=3 (2013)
P(t)=7t^2+8000
P(3)=7(3)^2+8000
P(3)=63+8000
P(3)=8063
.
t=6 (2016)
P(t)=7t^2+8000
P(6)=7(6)^2+8000
P(6)=252+8000
P(6)=8252
.
To plot, you simply plot the points on the graph:
(2010,8000)
(2011,8007)
(2012,8028)
(2013,8063)
(2016,8252)